# A nonequilibrium quantum Otto engine enhanced via multi-parameter control

### v1

@@ VII Tonks-Girardeau gas
Engine performance in the Tonks-Girardeau regime is demonstrated in Fig. 5, again in terms of net work and efficiency as a function of both quenched parameter ratios, $g_{h}/g_{l}$ and $\omega_{h}^{2}/\omega_{l}^{2}$. Here, our working fluid consists of N=20N\!=\!20 total atoms, at dimensionless interaction strength $\gamma_{0}\!\simeq\!8.5$ and dimensionless temperature $\tau_{0}\!=\!0.18$, meaning the system inhabits the low temperature regime of fermionization [67]. The Otto cycle investigated here is again thermo-chemical, where $\Delta N\!=\!2$ particles are exchanged with the reservoirs during the equilibration strokes, whereas the high and low equilibrium states are chosen to have a temperature ratio of $T_{h}/T_{l}\!=\!2$.
Notably, the enhancement to net work and efficiency achieved via the two-parameter cycle in the Tonks-Girardeau gas is not as significant when contrasted to operation in the quasicondensate regime. To compare the two-parameter engine cycle between these regimes, we consider the maximum net work normalized to the total particle number in the low energy equilibrium state, $N_{l}$, which varies significantly between the Tonks-Girardeau ($N_{l}\!=\!20$) and quasicondensate ($N_{l}\!=\!2000$) working fluids. Upon doing this, we find that the Otto cycle operating in the Tonks-Girardeau regime achieves a maximum net work of $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5\!\times\!10^{-2}$, which is two orders of magnitude less than the same cycle operating in the quasicondensate regime, where $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5$ (see Fig. 4 caption).
@@ VIII Conclusions
In this work, we investigated the operation of a sudden quench Otto cycle with control over multiple external parameters. To do this, we extended the work done recently in Ref. [21], for sudden quenches of single parameters, to the general case of arbitrary sets of controllable parameters. Under the sudden quench approximation, the total net work separates into a sum of its constituent quenches. From this, we derived general principles dictating when the net work extracted from a two-parameter sudden quench protocol exceeds that from related single-parameter cycles taken in isolation for arbitrary quantum models, benefiting an Otto engine performance. Further, this enhancement to the net work of the two-parameter Otto cycle was found to result in an improved coefficient of performance when operating as an Otto refrigerator.
The methods introduced were applied to the case of an Otto engine cycle in an experimentally realizable harmonically trapped 1D Bose gas with contact interactions. Control over both the strength of interactions and the harmonic trapping frequency was shown to result in a region of enhanced performance when both parameters are simultaneously quenched. This was analytically illustrated in the Thomas-Fermi approximation in the simplest example of a chemical engine operating in the weakly interacting regime of the 1D Bose gas at zero temperature.
The same enhancement to both net work and efficiency was demonstrated numerically using the thermodynamic Bethe ansatz in a finite temperature quasicondensate and the strongly interacting Tonks-Girardeau regime of the 1D Bose gas operating as a thermochemical engine. Finally, we highlight that the methods introduced here are applicable universally to quantum mechanical models, relying only on rapid control over multiple external parameters.
@@ Appendix A Sudden quench Otto cycle
Appendix B General multi-parameter enhancement for engine and refrigerator operation
Here, we examine an Otto cycle where the sudden quench is taken for the scalar strength parameters $c^{(1)}$ and $c^{(2)}$, as first introduced for a general physical model in Sec. II, and which are associated with the operators $\hat{\mathcal{V}}^{(1)}$ and $\hat{\mathcal{V}}^{(2)}$, respectively. The net work of this sudden quench Otto cycle is given by Eq. (4) of the main text, for a sudden quench over the parameters $c^{(1)}$ and $c^{(2)}$ between their respective values in the low ($l$) and high ($h$) energy equilibrium states.
To examine whether the net work of the two-parameter cycle exceeds the sum of the single-parameter cycles, we follow the same arguments laid out in the main text. In particular, we compare the net work of the two-parameter sub-cycle associated with $c^{(1)}$ with the single-parameter cycle associated with quenching only parameter $c^{(1)}$. The difference in net work for these contributions is given by
$-\Delta W^{(1)}\!\simeq$ $(c_{h}^{(1)}-c_{l}^{(1)})$
$\times$ $\left(\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{h}\right)}\!-\!\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{l}\right)}\right)$, (27)
which must be greater than zero for the two-parameter cycle to out-perform the single-parameter protocol. In particular, we require that $\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{h}\right)}\!>\!\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{l}\right)}$, i.e. we require the equilibrium expectation value of $\hat{\mathcal{V}}^{(1)}$ to increase with the complementary parameter, $c^{(2)}$. If this condition is met for both sub-cycles, then we arrive at the inequality
$\Delta W=-\Delta W^{(1)}-\Delta W^{(2)}>0,$ (28)
for the net work extracted from two-parameter Otto cycle to exceed the sum of the single-parameter cycles, where $\Delta W$ is defined in Eq. (13).
Finally, we note that, if the inequality $\Delta W\!>\!0$ is valid for a particular quantum model operating in an Otto cycle, it remains valid regardless of whether the Otto cycle operates as an engine, or one of the alternative protocols (e.g., refrigerator, thermal accelerator, or heater [19]). Indeed, considering the refrigerator operation as one of the alternative protocols [45, 19], shown schematically in Fig. 7, we first note that the key parameter is the coefficient of performance ($\mathrm{CoP}$), given by

### v2

@@ VI.2 Tonks-Girardeau gas
Engine performance in the Tonks-Girardeau regime is demonstrated in Fig. 4, again in terms of net work and efficiency as a function of both quenched parameter ratios, $g_{h}/g_{l}$ and $\omega_{h}^{2}/\omega_{l}^{2}$. Here, our working fluid consists of N=20N\!=\!20 total atoms, at dimensionless interaction strength $\gamma_{0}\!\simeq\!8.5$ and dimensionless temperature $\tau_{0}\!=\!0.18$, meaning the system inhabits the low temperature regime of fermionization [70]. The Otto engine investigated here is again thermo-chemical, where $\Delta N\!=\!2$ particles are exchanged with the reservoirs during the equilibration strokes, whereas the high and low equilibrium states are chosen to have a temperature ratio of $T_{h}/T_{l}\!=\!2$.
Notably, the enhancement to net work and efficiency achieved via the two-parameter cycle in the Tonks-Girardeau gas is not as significant when contrasted to operation in the quasicondensate regime. To compare the two-parameter engine cycle between these regimes, we consider the maximum net work normalized to the total particle number in the low energy equilibrium state, $N_{l}$, which varies significantly between the Tonks-Girardeau ($N_{l}\!=\!20$) and quasicondensate ($N_{l}\!=\!2000$) working fluids. Upon doing this, we find that the Otto engine operating in the Tonks-Girardeau regime achieves a maximum net work of $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5\!\times\!10^{-2}$, which is two orders of magnitude less than the same cycle operating in the quasicondensate regime, where $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5$ (see Fig. 3 caption).
@@
VII Transverse-field Ising model
As a final example of this general multi-parameter enhancement, we demonstrate its applicability in the context of the TFIM. This system consists of a 1D chain of spin-$1/2$ particles with nearest neighbor interactions in the presence of an external magnetic field. Unlike the harmonically trapped 1D Bose gas introduced in Sec. IV, the two-parameter Otto cycle for the TFIM has no uncontrolled Hamiltonian operators $\hat{H}_{0}$. Instead, we have first the external magnetic field $\mathcal{h}$ acting uniformly on the spins in the $z$-direction,
$c^{(1)}\hat{\mathcal{V}}^{(1)}=-\mathcal{h}\sum_{i=1}^{N}\hat{\sigma}^{z}_{i},$ (32)
with $\hat{\mathcal{V}}^{(1)}\equiv-\sum_{i=1}^{N}\hat{\sigma}^{z}_{i}$, where $\hat{\sigma}^{\alpha}_{i}$ ($\alpha=x,y,z$) are the Pauli matrices for the $i$-th lattice site, and $c^{(1)}\!\equiv\!\mathcal{h}$.
The only other term in the Hamiltonian for this model is given by the nearest-neighbour inter-site interaction in the $x$-direction, and is represented by the two-body operator,
$c^{(2)}\hat{\mathcal{V}}^{(1)}=-J\sum_{i=1}^{N}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1},$ (33)
with $\hat{\mathcal{V}}^{(2)}\equiv-\sum_{i=1}^{N}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}$, and $c^{(2)}\equiv J$ as the interaction strength. Here, we choose to work in the sector with both $\mathcal{h}\!>\!0$ and J>0J\!>\!0, though the analysis presented extends simply outside of this region.
Figure 5: Performance of the two-parameter sudden quench Otto engine cycle for the TFIM. All panels are presented in the same layout as in Fig. 3, adapted to the parameters utilized in the TFIM. Net work is given in natural units based on the low energy magnetic field strength, $\mathcal{h}_{l}$, and normalized to the total particle number. The low energy equilibrium state is fixed by $J_{l}/\mathcal{h}_{l}\!=\!1.2$, with $k_{B}T_{l}/\mathcal{h}_{l}\!=\!1$. The high energy equilibrium state has $T_{h}/T_{l}\!=\!2$. We observe a maximum net work of $W/N\mathcal{h}_{l}\!\simeq\!5\times 10^{-2}$, and an efficiency at maximum net work of $\eta\!\simeq\!0.3$.
The two-parameter sudden quench Otto engine cycle for the TFIM corresponds to a sudden quench over both the transverse magnetic field, $c^{(1)}\!=\!\mathcal{h}$, with $\mathcal{h}_{h}\!>\!\mathcal{h}_{l}$, and the interaction strength, $c^{(2)}\!=\!J$, with $J_{h}\!>\!J_{l}$. The net work of such an engine is then given by
$W$ $\!\simeq(\mathcal{h}_{h}\!-\!\mathcal{h}_{l})\left(\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle_{h}\!-\!\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle_{l}\right)$
$+\!(J_{h}\!-\!J_{l})\left(\sum_{i}\langle\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}\rangle_{h}\!-\!\sum_{i}\langle\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}\rangle_{l}\right).$ (34)
We focus here on the net work per particle, $W/N$, which may be calculated analytically at finite temperature in the thermodynamic limit via the Helmholtz free energy per particle, f=F/Nf\!=\!F/N. In particular, as shown in Ref. [81], in one dimension one may utilize the Jordan-Wigner transformation to map the spins to free fermions, which may be solved exactly to give,
$f\!=\!-\frac{1}{k_{B}T}\left\{\ln 2\!+\!\frac{1}{\pi}\int_{0}^{\pi}dk\ln\left[\cosh\left(\frac{\Lambda_{k}}{k_{B}T}\right)\right]\right\},$ (35)
where $\Lambda_{k}=\sqrt{J^{2}+\mathcal{h}^{2}-2J\mathcal{h}\cos k}$ depends on both the strength of the transverse magnetic field, $\mathcal{h}$, and on the inter-site interaction strength, $J$.
The transverse magnetization per particle, $m_{z}\!=\!\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle/N$—appearing on the first line of Eq. (34)—may then be evaluated using the Hellmann-Feynman theorem [52, 53],
$m_{z}=-\frac{\partial f}{\partial\mathcal{h}}=\frac{1}{\pi}\int_{0}^{\pi}dk\frac{\mathcal{h}-J\cos k}{\Lambda_{k}}\tanh\left(\frac{\Lambda_{k}}{k_{B}T}\right).$ (36)
One may utilize this method also to calculate the nearest-neighbor correlation per particle, $\sum_{i}\langle\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}\rangle/N$, which is required for evaluating the remaining terms in Eq. (34); this takes the same form as the transverse net magnetization above, only swapping $\mathcal{h}\leftrightarrow J$ due to the symmetry of the Helmholtz free energy in these parameters.
These analytically calculable operator expectation values may then be used to evaluate the performance of the two-parameter Otto engine cycle for the TFIM in the thermodynamic limit. In Figs.  5 (a) and (b), we show the net work per particle and the efficiency of such an engine cycle. As we see from these figures, the net work per particle is low, being of the same order of magnitude achieved in the Tonks-Girardeau regime of the 1D Bose gas shown in Fig. 4. The efficiency, however, is significantly higher than that achieved from a multi-parameter quench in the 1D Bose gas.
The relative increase to efficiency for the Otto engine cycle in the TFIM, when contrasted with that obtained from the 1D Bose gas, results from the fact that there is no explicit kinetic energy term present in the Hamiltonian of the TFIM. The thermal energy absorbed by the system from the hot reservoir is therefore exclusively transferred into the controllable degrees of freedom utilized in the Otto engine cycle, with no losses to the kinetic energy term, $\hat{H}_{0}$, which is absent in the TFIM.
Multi-parameter enhancement to both the net work and efficiency over the single-parameter engine cycles are shown in Figs. 5(c) and (d), respectively. Here, as was the case for the 1D Bose gas, we observe a significant enhancement in both performance measures. In particular, we observe that the maximum net work per particle arising from the multi-parameter quench significantly out-performs the sum of that from its single-parameter counterparts.
VII.1 Small parameter quenches
Multi-parameter enhancement to the net work for small parameter quench sizes was investigated in Sec. III.1. There, we found a general inequality for $\Delta W$ in terms of the second mixed derivative of the Helmholtz free energy, given in Eq. (19). The TFIM possesses a quantum critical region at finite temperature extending from the zero-temperature critical point $\mathcal{h}\!=\!J$ [82], which directly affects the second mixed derivative of the Helmholtz free energy, otherwise known as the static susceptibility. Here, we investigate multi-parameter enhancement to net work under small parameter quenches within this critical region.
We begin by defining the infinitesimal quench of the transverse field, $\mathcal{h}$, and inter-site interaction, $J$, as $d\mathcal{h}\!=\!\mathcal{h}_{h}\!-\!\mathcal{h}_{l}\!>\!0$ and $dJ\!\equiv\!J_{h}\!-\!J_{l}\!>\!0$, respectively. Then, as we are considering operation in the thermodynamic limit, we express the condition $\Delta W\!>\!0$, given in Eq. (19), in terms of the net work enhancement per particle for a two-parameter Otto engine,
$\frac{-\Delta W}{N}=2\,d\mathcal{h}\,dJ\frac{\partial^{2}f}{\partial\mathcal{h}\partial J}>0,$ (37)
where f=F/Nf\!=\!F/N is the Helmholtz free energy per particle as before. As noted in Sec. III.1, multi-parameter enhancement to the net work for small parameter quenches is only dependent on the properties of the high energy equilibrium state. As such, in the following analysis we omit (for clarity) the subscripts defining the high and low energy equilibrium states, instead noting here that all parameters are taken at their high energy equilibrium values.
Using Eq. (35) for $f$, we may evaluate its mixed second derivative,
$\frac{\partial^{2}f}{\partial\mathcal{h}\partial J}$ $=\frac{1}{\pi}\int_{0}^{\pi}\bigg[\frac{\mathcal{h}J\sin^{2}(k)}{\Lambda_{k}^{3}}\tanh\left(\frac{\Lambda_{k}}{k_{B}T}\right)$
$-\!\frac{(\mathcal{h}\!-\!J\cos k)(J\!-\!\mathcal{h}\cos k)}{k_{B}T\Lambda_{k}^{3}}\mathrm{sech}^{2}\left(\frac{\Lambda_{k}}{k_{B}T}\right)\bigg].$ (38)
In Fig. 6, we show the behavior of this static susceptibility for three distinct values of the temperature. Notably, as temperature decreases, the static susceptibility shows a smooth peak at the critical line $J\!=\!\mathcal{h}$. This arises from the first term in Eq. (VII.1), which remains non-zero in the zero-temperature limit and is known to diverge at T=0T\!=\!0 for $J\!=\!\mathcal{h}$. According to Eq. (37), the presence of this peak implies an increase to the two-parameter enhancement of the net work, implying that operation of this Otto cycle is optimal at the critical line for the case of small parameter quenches.
Figure 6: The mixed second derivative of the free energy, otherwise known as the static susceptibility, for the TFIM. Two-parameter enhancement to the net work is directly proportional to the static susceptibility for small quenches (see text). Under decreasing temperature, a smooth peak emerges at the critical line, $J\!=\!\mathcal{h}$, indicating an increase to the two-parameter enhancement for the net work over the sum of the single-parameter sub-cycles.
Though the enhancement to net work decreases with increasing temperature, it remains positive at arbitrarily high temperatures. To demonstrate this, we Taylor expand the mixed second derivative, given in Eq. (VII.1), for $k_{B}T\!\gg\!\mathcal{h},J$,
$\frac{\partial^{2}f}{\partial\mathcal{h}\partial J}\simeq\frac{4\mathcal{h}J}{3(k_{B}T)^{3}}>0.$ (39)
The static susceptibility is therefore strictly positive for the parameter regime we are investigating, i.e. for $\mathcal{h}\!>\!0$ and J>0J\!>\!0, indicating a two-parameter enhancement to the net work at arbitrarily high temperatures.
The analysis presented in this section highlights the utility of small multi-parameter quenches for models which possess analytic expressions for their Helmholtz free energy. The simplest example of such models are free bosons or free fermions, and models which map onto these, such as the XY model, which is a generalization of the TFIM investigated here [83]. Additionally, one may consider Bethe ansatz integrable models, such as the Yang-Gaudin model for spin-$1/2$ particles in 1D, or the Heisenberg spin chain [83], for which there are again analytic expressions for the free energy. Notably, the Heisenberg XYZ model possesses three individually tunable parameters, and would therefore present an interesting extension to this work beyond the two-parameter case [84, 83].
@@ VIII Conclusions
In this work, we investigated the operation of a sudden quench Otto cycle with control over multiple external parameters. To do this, we extended the work done recently in Ref. [21], for sudden quenches of single parameters, to the general case of arbitrary sets of controllable parameters. Under the sudden quench approximation, the total net work separates into a sum of its constituent quenches. From this, we derived general principles dictating when the net work extracted from a two-parameter sudden quench protocol exceeds that from related single-parameter cycles taken in isolation for arbitrary quantum models, benefiting Otto engine performance. Further, this enhancement to the net work of the two-parameter Otto engine was found to result in an improved coefficient of performance when operating as an Otto refrigerator.
The methods introduced were applied to the case of an Otto engine cycle in an experimentally realizable harmonically trapped 1D Bose gas with contact interactions. Control over both the strength of interactions and the harmonic trapping frequency was shown to result in a region of enhanced performance when both parameters are simultaneously quenched. This was analytically illustrated in the Thomas-Fermi approximation in the simplest example of a chemical engine operating in the weakly interacting regime of the 1D Bose gas at zero temperature. The same enhancement to both net work and efficiency was demonstrated numerically using the thermodynamic Bethe ansatz in a finite temperature quasicondensate and the strongly interacting Tonks-Girardeau regime of the 1D Bose gas operating as a thermochemical engine.
To emphasize the universality of the methods introduced in this work to arbitrary quantum mechanical models, we applied the multi-parameter Otto cycle to the transverse-field Ising model, where we again observed a general two-parameter enhancement to both net work and efficiency. Finally, for small quench sizes of multiple parameters, we derived a general inequality for the multi-parameter enhancement of the net work based on the Helmholtz free energy. This was then used to provide analytic insight into the operation of the two-parameter sudden quench Otto engine when applied to the transverse-field Ising model.
We emphasize that the analysis presented in this work is restricted to the sudden-quench regime, where the work depends only on the initial expectation values of the relevant operators and is therefore independent of the path taken in parameter space. For finite-time protocols, the work generally becomes path-dependent due to nonadiabatic excitations and the dynamical evolution of correlations. While the enhancement mechanism identified here is expected to persist qualitatively for sufficiently fast protocols, a quantitative description of general finite-time paths would require solving the full nonequilibrium dynamics and is beyond the scope of the present work. We also note that shortcuts to adiabaticity [85] are available individually for control over both interaction strengths [27] and harmonic trapping frequency [86, 87, 88], for the 1D Bose gas. It would therefore be an interesting question as to whether these shortcut methods can be combined to effectively enhance the performance of the Otto engine under a simultaneous two-parameter control.
@@ Appendix A Sudden quench Otto cycle
Appendix B Refrigerator enhancement
If the inequality $\Delta W\!>\!0$ is valid for a particular quantum model operating in an Otto cycle, it remains valid regardless of whether the Otto cycle operates as an engine, or one of the alternative protocols (e.g., refrigerator, thermal accelerator, or heater [19]). Indeed, considering the refrigerator operation as one of the alternative protocols [45, 19], shown schematically in Fig. 8, we first note that the key parameter is the coefficient of performance ($\mathrm{CoP}$), given by